Saskatchewan River Enterprises​ (SRE) has

​$700700

million in debt and

1010

million shares of equity outstanding. Its excess cash reserves are

​$2525

million. SRE is expected to generate

​$300300

million in free cash flows next year with a growth rate of​ 2% per year in perpetuity.​ SRE's cost of equity capital is

1212​%.

a. What is​ SRE's stock​ price?

b. After analyzing the​ company, you believe that the growth rate should be​ 3% instead of​ 2%. Assume debt and cash values do not change. What should the stock price​ be, given the higher growth​ rate?

c. Given the growth rate change from​ 2% to​ 3%, how can you calculate the change in stock price without calculating the amounts in a or b​ above?
Answer #1

a) As per earnings growth model:

Value of Business = Value of future cash flows (1+g)/ ke-g

g = growth rate=2%

ke = Cost of equity=12.%

Value of business= $ 300(1+0.02)/0.12-002

=$ 3060 million

Value of share = Value of business/number of share

= 3060/25=$122.4 per share

b)  

Value of Business = Value of future cash flows (1+g)/ ke-g

g = growth rate=3%

ke = Cost of equity=12.%

Value of business= $ 300(1+0.03)/0.12-003

=$ 3400 million

Value of share = Value of business/number of share

= 3400/25=$136 per share